OpenStudy (anonymous):

PLEASE HELP! I WILL FAN AND MEDAL!!!!!!!! What is the ratio of X to Y? 3 2X - = ----- 2 2X+Y A. -3/2 B. -2/3 C. 2/3 D. 3/2

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

@shaik0124

OpenStudy (shaik0124):

cross multiply

OpenStudy (anonymous):

c?

OpenStudy (shaik0124):

6x+3y=4x 6x-4x=-3y 2x=-3y x/y=-3/2

OpenStudy (anonymous):

ohhh ok I get it know! THank you!